OpenStudy (anonymous):

The graph below plots a function f(x): graph of line segment going through ordered pairs 0, 100 and 3, 220 If x represent time, the average rate of change of the function f(x) in the first three seconds is ___. Numerical Answers Expected!

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

Average rate of change it just the slope

OpenStudy (anonymous):

\[\huge m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Your ordered pairs are (0,100) and (3,220) just plug and chug

OpenStudy (anonymous):

?

OpenStudy (anonymous):

You have to tell me what part you don't understand giving me a question mark won't help.

OpenStudy (anonymous):

The points you're given

OpenStudy (anonymous):

@iambatman i do not understand who to plyg and chug xd

OpenStudy (anonymous):

(0,100) and (3,220)

OpenStudy (anonymous):

where do plug those into the equation though

OpenStudy (anonymous):

The first coordinate is (x1,y1) and the other is (x2,y2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x1=0 y1,= 100 x2=3 y2=220 @iambatman like that?

OpenStudy (anonymous):

Yeah, now use the slope formula

OpenStudy (anonymous):

The slope formula is what I wrote above, did you get it?

OpenStudy (anonymous):

no @iambatman

OpenStudy (anonymous):

Well show me an attempt and I'll tell you if you're right or not

OpenStudy (anonymous):

m=2202-1001/32-01 @iambatman

OpenStudy (anonymous):

2202? 1001?

OpenStudy (anonymous):

\[m=\frac{ 220-100 }{ 3-0 }\]

OpenStudy (anonymous):

yes @iambatman

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

220-100/3 =?

OpenStudy (anonymous):

33.333333333... @iambatman

OpenStudy (anonymous):

What's 220 - 100?

OpenStudy (anonymous):

120

OpenStudy (anonymous):

the answer is 40 @iambatman

OpenStudy (anonymous):

Yes